# DM-Codec: Distilling Multimodal Representations for Speech Tokenization

### v1

@@ 6 Limitations and Broader Impact
Broader Impact. The integration of language models in speech processing has traditionally focused on model-specific implementations or specific training objectives. In this work, we propose a novel approach by leveraging a language model during the tokenization phase through our model, DM-Codec. By incorporating language-specific representations from the corresponding text, DM-Codec enhances the quality of discrete speech representations. This method bridges the gap between language and speech models, offering a more unified approach to multimodal representation learning. DM-Codec provides a robust framework for generating high-quality audio representations, with potential applications in various domains, including multilingual speech processing, low-resource languages, and other audio-related tasks. Our findings pave the way for more effective and contextually aware speech processing models, contributing to advancements in the broader field of speech and language technologies.
7 Conclusion
In this work, we introduced a speech tokenizer DM-Codec, with two novel distillation methods to leverage multimodal (acoustic, semantic, and contextual) representations from a language model and speech self-supervised learning model. Our extensive experimental results and ablation studies suggest that distilling multimodal representations enables DM-Codec to introduce salient speech information in discrete speech tokens. Our significance analysis further revealed that DM-Codec with comprehensive multimodal representations consistently outperforms existing speech tokenizers. This approach highlights the potential of multimodal representations to enhance speech tokenization in various domains, including multilingual and code-switched speech processing.

### v2

@@ Appendix G Broader Impact and Potential Risks
The integration of language models in speech processing has traditionally focused on model-specific implementations or specific training objectives. In this work, we propose a novel approach by leveraging language models during the tokenization phase through our model, DM-Codec. By incorporating language-specific representations from the corresponding text, DM-Codec enhances the quality of discrete speech representations. This method bridges the gap between language and speech models, offering a more unified approach to multimodal representation learning. DM-Codec provides a robust framework for generating high-quality audio representations, with potential applications in various domains, including multilingual speech processing, low-resource languages, and other audio-related tasks. Our findings pave the way for more effective and contextually aware speech processing models, contributing to advancements in the broader field of speech and language technologies. Despite its advancements, DM-Codec-TTS poses risks of misuse, such as voice spoofing or impersonation, due to its ability to synthesize speech while preserving speaker identity. Our future work will prioritize robust detection mechanisms and ethical safeguards to ensure responsible use.
Appendix H Reconstructed Speech Comparison
Figure 6 compares the Mel-Spectrograms of the original speech with reconstructed speech from DM-Codec, EnCodec, SpeechTokenizer, and FACodec.
Refer to caption Refer to caption
(a) Original Speech 1
Refer to caption Refer to caption
(b) Original Speech 2
Refer to caption Refer to caption
(c) DM-Codec Speech 1
Refer to caption Refer to caption
(d) DM-Codec Speech 2
Refer to caption Refer to caption
(e) EnCodec Speech 1
Refer to caption Refer to caption
(f) EnCodec Speech 2
Refer to caption Refer to caption
(g) SpeechTokenizer Speech 1
Refer to caption Refer to caption
(h) SpeechTokenizer Speech 2
Refer to caption Refer to caption
(i) FACodec Speech 1
Refer to caption Refer to caption
(j) FACodec Speech 2
Figure 6: Reconstructed speech examples with clickable play buttons above each Mel-spectrogram.